# One protein is all you need

### v2

@@ 1 Introduction
Figure 1: Example of protein structure prediction after single-protein model customization via ProteinTTT. ESMFold poorly predicts the structure of the CASP14 target T1074 (white) because the underlying language model ESM2 poorly fits the sequence, as indicated by the high perplexity (left and Fig. 2E in Lin et al. (2023)). Self-supervised test-time customization of ESM2 to the single sequence of T1074 reduces the perplexity, resulting in improved structure prediction (right).
A comprehensive understanding of protein structure, function, and fitness is essential for advancing research in the life sciences (Subramaniam & Kleywegt, 2022; Tyers & Mann, 2003; Papkou et al., 2023). While machine learning models have shown remarkable potential in protein research, they are typically optimized for achieving the best average performance across large datasets (Jumper et al., 2021; Watson et al., 2023; Kouba et al., 2023). However, biologists often focus their research on individual proteins or protein complexes involved in, for example, metabolic disorders (Ashcroft et al., 2023; Gunn & Neher, 2023), oncogenic signaling (Hoxhaj & Manning, 2020; Keckesova et al., 2017), neurodegeneration (Gulen et al., 2023; oh Seo et al., 2023), and other biological phenomena (Gu et al., 2022). In these scenarios, detailed insights into a single protein can lead to significant scientific advances.
However, general machine learning models for proteins often struggle to generalize to practically interesting individual cases due to data scarcity (Bushuiev et al., 2023; Chen & Gong, 2022) and distribution shifts (Škrinjar et al., 2025; Tagasovska et al., 2024; Feng et al., 2024). Bridging the gap between broad, dataset-wide optimization and precision needed to study single proteins of practical interest remains a key challenge in integrating machine learning into biological research (Sapoval et al., 2022). This challenge is particularly acute in computational biology, where accurate predictions for individual proteins are essential to guide resource-intensive wet-lab experiments, in contrast to domains such as natural language processing or computer vision, where models are typically expected to flexibly handle diverse prompts from many users in real time (Brown, 2020; Ramesh et al., 2021).
@@ 2 Background and related work
In machine learning applied to proteins, architectures often follow a Y-shaped paradigm (Gandelsman et al., 2022), consisting of a backbone feature extractor $f$ operating on protein tokens $x$, a self-supervised head $g$, and an alternative fine-tuning head $h$. During training, $g\circ f$ is first pre-trained, and the pre-trained backbone $f$ is then reused to fine-tune $h\circ f$ toward a downstream task. Here, $\circ$ denotes a composition of two machine learning modules (e.g., $g$ is applied on top of $f$ in $g\circ f$). At test time, the final model $h\circ f$ is fixed. Generalization is achieved by leveraging the rich knowledge encoded in the backbone $f$ and the task-specific priors embedded in the fine-tuning head $h$. This paradigm enables overcoming data scarcity during fine-tuning and underlies breakthrough approaches in protein structure prediction (Lin et al., 2023), protein design (Watson et al., 2023), protein function prediction (Yu et al., 2023), and other tasks (Hayes et al., 2024).
The backbone $f$ is typically a large neural network pre-trained in a self-supervised way on a large dataset using a smaller pre-training projection head $g$ (Hayes et al., 2024). The fine-tuning head $h$, however, depends on the application. In some cases, $h$ is a large neural network, repurposing the pre-trained model entirely (Watson et al., 2023; Lin et al., 2023); in others, $h$ is a minimal projection with few parameters (Cheng et al., 2023), or even without any parameters at all (i.e., a zero-shot setup; Meier et al. (2021); Dutton et al. (2024)). The fine-tuning head $h$ can also be a machine learning algorithm other than a neural network (Samusevich et al., 2025).
@@
While the objective of fine-tuning $h\circ f$ is determined by the downstream application, the choice of pre-training objective for $g\circ f$ is less straightforward. Nevertheless, the dominant paradigm for protein pre-training is masked modeling, which optimizes model weights to reconstruct missing protein parts. This objective has proven effective across diverse tasks (Heinzinger & Rost, 2025; Schmirler et al., 2024), including structure (Lin et al., 2023; Jumper et al., 2021), fitness (Meier et al., 2021; Su et al., 2023), and function prediction (Samusevich et al., 2025; Yu et al., 2023; Elnaggar et al., 2021), as well as protein design (Hayes et al., 2024), and has been successfully applied to various protein representations such as sequences (Hayes et al., 2024; Elnaggar et al., 2023), graphs (Dieckhaus et al., 2024; Bushuiev et al., 2023), and voxels (Diaz et al., 2023).
@@
Several studies have shown that machine learning models for proteins benefit from being fine-tuned on protein-specific  (Notin et al., 2024; Kirjner et al., 2023; Rao et al., 2019) or protein family-specific (Sevgen et al., 2023; Samusevich et al., 2025) data. However, collecting additional data may be resource-intensive, and for many targets, relevant datasets or proteins may be limited or not available (Durairaj et al., 2023; Kim et al., 2025). In this paper, we propose a versatile method enabling customizing PLMs for a single target protein or its MSA in a self-supervised manner, on the fly, and without assuming any additional data. Customization methods have been developed in computer vision (Chi et al., 2024; Wang et al., 2023; Xiao et al., 2022; Karani et al., 2021) and natural language processing (Hübotter et al., 2024; Hardt & Sun, 2023; Ben-David et al., 2022; Banerjee et al., 2021). The paradigm of test-time training (TTT), developed to mitigate distribution shifts in computer vision applications (Gandelsman et al., 2022; Sun et al., 2020), is the main inspiration for our work. We demonstrate that customization via test-time training enhances the accuracy of PLMs across a wide range of downstream tasks even without the presence of explicit distribution shifts.
@@ 3.1 Self-supervised customization to a target protein
where $\theta_{0}$ denotes pre-trained parameters and $\theta_{x}$ parameters optimized for the target protein $x$ using the self-supervised track $g\circ f$, while the supervised head $h$ remains frozen. Figure˜2a illustrates our customization approach, which is summarized in the following sections. Appendix˜B describes the extension of our method to customization using a MSA of a protein, rather than its single sequence.
@@
We customize $g\circ f$ to a single target protein sequence $x$ via minimizing the masked language modeling objective (Devlin, 2018; Rives et al., 2021):
@@
where $x$ denotes a sequence of protein tokens (typically amino acid types), and $\mathbb{E}_{M}$ represents the expectation over randomly sampled masking positions $M$. The objective function $\mathcal{L}(x;\theta)$ maximizes the log-probabilities $\log{p(x_{i}|x_{\setminus M};\theta)}\ \dot{=}\ g(f(x_{\setminus M};\theta))_{i}$ of the true (i.e., wild-type) tokens $x_{i}$ at the masked positions $i\in M$ in the partially masked sequence $x_{\setminus M}$, where $\theta$ denotes the parameters of the backbone $f$, and $g$ is the masked language modeling head. Please note that here we focus on bi-directional masked modeling models, which employ random masking, but the method can be easily extended to models employing autoregressive masking.
To ensure consistency between the customization and pre-training, ProteinTTT adopts the same masking and data preprocessing strategies used during pre-training. Specifically, $p_{\text{mask}}(M)$ can follow different distributions, such as sampling a fixed proportion (e.g., 15%) of random amino acid tokens (Lin et al., 2023), or dynamically varying the number of sampled tokens based on another distribution (e.g., a beta distribution; Hayes et al. (2024)). During the customization, we replicate the masking distribution used during the pre-training. We also replicate other pre-training practices, such as replacing 10% of masked tokens with random tokens and another 10% with the original tokens (Devlin, 2018; Lin et al., 2023; Su et al., 2023) or cropping sequences to random 1024-token fragments (Lin et al., 2023; Su et al., 2023).
@@
Since customization with ProteinTTT does not assume more than a single protein, early stopping on validation data is not feasible. To address this, we first fine-tune the pre-trained parameters $\theta_{0}$ of a backbone $f$ for a fixed number of steps $T$, yielding parameters $\Theta=\{\theta_{0},\theta_{1},\dots,\theta_{T}\}$. The final customized parameters $\theta_{x}$ are selected as $\operatorname*{arg\,max}_{\theta\in\Theta}c(h(f(x;\theta)))$ where $c$ is a confidence function. If $c$ is not available, we set $\theta_{x}=\theta_{T}$. Section˜G.2 discusses how using pLDDT as the confidence function $c$ for structure prediction makes ProteinTTT robust to hyperparameter selection and how the number of steps $T$ can be fixed (e.g., $T=30$) while optimizing learning rate and batch size effectively. Before customizing for the next target protein, the parameters are reset to $\theta_{0}$.
To make ProteinTTT easily applicable to large-scale models (e.g., the 3B-parameter ESM2 backbone), we leverage low-rank adaptation (LoRA; Hu et al. (2021)) and gradient accumulation during customization. Additionally, to improve the stability and predictability of customization, we use stochastic gradient descent (SGD; Ruder (2016)) instead of the commonly used Adam optimizer (Kingma & Ba, 2015), following (Gandelsman et al., 2022). Further details are provided in Appendix˜E.
@@ 3.2 Inference on downstream tasks
Since Y-shaped architectures are prevalent in protein machine learning, ProteinTTT can be straightforwardly applied to numerous tasks. In this work, we consider three standard problems: protein structure, fitness, and function prediction, and apply our method to corresponding well-established models. For structure prediction, we apply ProteinTTT to ESMFold (Figure˜3, Lin et al. (2023), HelixFold-Single (Fang et al., 2023), and ESM3 (Hayes et al., 2024); for fitness prediction, we use ESM2 (Lin et al., 2023), SaProt (Su et al., 2023), ProSST (Li et al., 2024), and MSA Transformer (Rao et al., 2021); and for function prediction, we apply ProteinTTT to ESM-1v-based (Meier et al., 2021) EnzymeExplorer (Samusevich et al., 2025) and ESM-1b-based (Rives et al., 2021) Light attention (Stärk et al., 2021).
In all models we consider, $f$ is a Transformer encoder that takes protein tokens as input, and $g$ is a masked language modeling head (a layer mapping token embeddings to amino acid types). The downstream task heads $h$ vary strongly across tasks. For structure prediction, $h$ is a protein structure predictor: in ESMFold and HelixFold-Single, it is an AlphaFold2-inspired module (Jumper et al., 2021), while in ESM3, it is a VQ-VAE structure decoder (Razavi et al., 2019). For fitness prediction, $h$ outputs a single score per sequence; ESM2, SaProt, and ProSST perform zero-shot inference using $h\circ f$ via log odds from $g$, with $h$ functioning as a simple adaptation of $g$ without introducing extra parameters. The function predictors are classification models: in EnzymeExplorer (Samusevich et al., 2025), $h$ is a random forest that outputs substrate probabilities, and in Light attention (Stärk et al., 2021), $h$ is a light attention module predicting protein localization classes within a cell.
@@ 4.1 Protein structure prediction
Protein structure prediction is the task of predicting 3D coordinates of protein atoms given the amino acid sequence. It is arguably one of the best-established problems in computational biology (Jumper et al., 2021; Lin et al., 2023; Abramson et al., 2024).
@@
To evaluate the performance of ProteinTTT, we employ CAMEO, a standard benchmark for protein folding. We use the validation and test folds from Lin et al. (2023), focusing only on targets with low-confidence predictions from the base ESMFold, as determined by pLDDT and perplexity (Section˜E.1). We use the standard TM-score (Zhang & Skolnick, 2004) and LDDT (Mariani et al., 2013) metrics to evaluate global and local structure prediction quality, respectively.
As baseline methods, we use techniques alternative to ProteinTTT for improving the performance of the pre-trained base models. In particular, the ESMFold paper proposes randomly masking 15% of amino acids in a protein sequence before inference, allowing for sampling multiple protein structure predictions from the regression ESMFold model (Lin et al., 2023). For each sequence, we sample a number of predictions equal to the total number of ProteinTTT steps and refer to this baseline as ESMFold + MP (Masked Prediction). As a baseline for ESM3, we use chain-of-thought iterative decoding, referred to as ESM3 + CoT, proposed in the ESM3 paper (Hayes et al., 2024).
@@ Results.
Customization with ProteinTTT consistently improves the performance of all the tested methods, ESMFold, HelixFold-Single, and ESM3, outperforming the masked prediction (ESMFold + MP) and chain-of-thought (ESM3 + CoT) baselines, as shown in Table˜2. Among the 18 challenging CAMEO test proteins, ProteinTTT significantly improved the prediction of 7, 5, and 6 structures from ESMFold, HelixFold-Single, and ESM3, respectively, while only slightly disrupting the prediction of 2, 1, and 1 structures, respectively (Figure˜A6). Most notably, ProteinTTT enables accurate structure prediction for targets that are poorly predicted with the original models. For instance, Figure˜1 presents a strongly improved structure predicted using ESMFold + ProteinTTT for the target that was part of the CASP14 competition and shown as an unsuccessful case in the original ESMFold publication (Lin et al. (2023), Fig. 2E). Another example is shown in Figure˜3, where ProteinTTT refined the structure prediction from a low-quality prediction (TM-score =0.29=0.29) to a nearly perfectly folded protein (TM-score =0.92=0.92). Figure˜A4 shows that ESMFold + ProteinTTT maintains computational efficiency of ESMFold, being an order of magnitude faster than AlphaFold2. Figure˜A11 additionally demonstrates the robustness of ESM3 + ProteinTTT to the choice of hyperparameters.

### v3

@@ 1 Introduction
Figure 1: Example of protein structure prediction after single-protein model customization via ProteinTTT. ESMFold poorly predicts the structure of the CASP14 target T1074 (white) because the underlying language model ESM2 poorly fits the sequence, as indicated by the high perplexity (left and Fig. 2E in lin2023evolutionary). Self-supervised test-time customization of ESM2 to the single sequence of T1074 reduces the perplexity, resulting in improved structure prediction (right).
A comprehensive understanding of protein structure, function, and fitness is essential for advancing research in the life sciences (subramaniam2022paradigm; tyers2003from; papkou2023rugged). While machine learning models have shown remarkable potential in protein research, they are typically optimized for achieving the best average performance across large datasets (jumper2021highly; watson2023novo; kouba2023machine). However, biologists often focus their research on individual proteins or protein complexes involved in, for example, metabolic disorders (ashcroft2023glucokinase; gunn2023structure), oncogenic signaling (hoxhaj2020pi3k; keckesova2017lactb), neurodegeneration (gulen2023cgas; seo2023apoe), and other biological phenomena (gu2022sestrin). In these scenarios, detailed insights into a single protein can lead to significant scientific advances.
However, general machine learning models for proteins often struggle to generalize to practically interesting individual cases due to data scarcity (bushuiev2023learning; chen2022hotprotein) and distribution shifts (vskrinjar2025have; tagasovska2024antibody; feng2024deep). Bridging the gap between broad, dataset-wide optimization and precision needed to study single proteins of practical interest remains a key challenge in integrating machine learning into biological research (sapoval2022current). This challenge is particularly acute in computational biology, where accurate predictions for individual proteins are essential to guide resource-intensive wet-lab experiments, in contrast to domains such as natural language processing or computer vision, where models are typically expected to flexibly handle diverse prompts from many users in real time (brown2020language; ramesh2021zero).
@@ 2 Background and related work
In machine learning applied to proteins, architectures often follow a Y-shaped paradigm (gandelsman2022test), consisting of a backbone feature extractor $f$ operating on protein tokens $x$, a self-supervised head $g$, and an alternative fine-tuning head $h$. During training, $g\circ f$ is first pre-trained, and the pre-trained backbone $f$ is then reused to fine-tune $h\circ f$ toward a downstream task. Here, $\circ$ denotes a composition of two machine learning modules (e.g., $g$ is applied on top of $f$ in $g\circ f$). At test time, the final model $h\circ f$ is fixed. Generalization is achieved by leveraging the rich knowledge encoded in the backbone $f$ and the task-specific priors embedded in the fine-tuning head $h$. This paradigm enables overcoming data scarcity during fine-tuning and underlies breakthrough approaches in protein structure prediction (lin2023evolutionary), protein design (watson2023novo), protein function prediction (yu2023enzyme), and other tasks (hayes2024simulating).
The backbone $f$ is typically a large neural network pre-trained in a self-supervised way on a large dataset using a smaller pre-training projection head $g$ (hayes2024simulating). The fine-tuning head $h$, however, depends on the application. In some cases, $h$ is a large neural network, repurposing the pre-trained model entirely (watson2023novo; lin2023evolutionary); in others, $h$ is a minimal projection with few parameters (cheng2023accurate), or even without any parameters at all (i.e., a zero-shot setup; meier2021language; dutton2024improving). The fine-tuning head $h$ can also be a machine learning algorithm other than a neural network (samusevich2024discovery).
@@
While the objective of fine-tuning $h\circ f$ is determined by the downstream application, the choice of pre-training objective for $g\circ f$ is less straightforward. Nevertheless, the dominant paradigm for protein pre-training is masked modeling, which optimizes model weights to reconstruct missing protein parts. This objective has proven effective across diverse tasks (heinzinger2025teaching; schmirler2024fine), including structure (lin2023evolutionary; jumper2021highly), fitness (meier2021language; su2023saprot), and function prediction (samusevich2024discovery; yu2023enzyme; elnaggar2021prottrans), as well as protein design (hsieh2025elucidating; hayes2024simulating; nijkamp2023progen2), and has been successfully applied to various protein representations such as sequences (hayes2024simulating; elnaggar2023ankh), graphs (dieckhaus2024transfer; bushuiev2023learning), and voxels (diaz2023stability).
@@
Several studies have shown that machine learning models for proteins benefit from being fine-tuned on protein-specific  (notin2024proteingym; kirjner2023improving; rao2019evaluating) or protein family-specific (sevgen2023prot; samusevich2024discovery) data. However, collecting additional data may be resource-intensive, and for many targets, relevant datasets or proteins may be limited or not available (durairaj2023uncovering; kim2025bfvd). In this paper, we propose a versatile method enabling customizing PLMs for a single target protein or its MSA in a self-supervised manner, on the fly, and without assuming any additional data. Customization methods have been developed in computer vision (chi2024adapting; wang2023test; xiao2022learning; karani2021test) and natural language processing (hubotter2024efficiently; hardt2023test; ben2022pada; banerjee2021self). The paradigm of test-time training (TTT), developed to mitigate distribution shifts in computer vision applications (gandelsman2022test; sun2020test), is the main inspiration for our work. We demonstrate that customization via test-time training enhances the accuracy of PLMs across a wide range of downstream tasks even without the presence of explicit distribution shifts.
@@ 3.1 Self-supervised customization to a target protein
where $\theta_{0}$ denotes pre-trained parameters and $\theta_{x}$ parameters optimized for the target protein $x$ using the self-supervised track $g\circ f$, while the supervised head $h$ remains frozen. Figure˜2a illustrates our customization approach, which is summarized in the following sections. Appendix˜C describes the extension of our method to customization using a MSA of a protein, rather than its single sequence.
@@
We customize $g\circ f$ to a single target protein sequence $x$ via minimizing the masked language modeling objective (devlin2018bert; rives2021biological):
@@
where $x$ denotes a sequence of protein tokens (typically amino acid types), and $\mathbb{E}_{M}$ represents the expectation over randomly sampled masking positions $M$. The objective function $\mathcal{L}(x;\theta)$ maximizes the log-probabilities $\log{p(x_{i}|x_{\setminus M};\theta)}\ \dot{=}\ g(f(x_{\setminus M};\theta))_{i}$ of the true (i.e., wild-type) tokens $x_{i}$ at the masked positions $i\in M$ in the partially masked sequence $x_{\setminus M}$, where $\theta$ denotes the parameters of the backbone $f$, and $g$ is the masked language modeling head. While we focus on classical bi-directional masked modeling, we also demonstrate that ProteinTTT can be similarly applied to autoregressive and discrete diffusion models (Appendix˜B).
To ensure consistency between customization and pre-training, ProteinTTT adopts the same masking and preprocessing strategies used during pre-training. Specifically, $p_{\text{mask}}(M)$ can follow different distributions, such as sampling a fixed ratio (e.g., 15%) of random tokens (lin2023evolutionary), or dynamically varying the number of sampled tokens based on another distribution (e.g., a beta distribution; hayes2024simulating). During customization, we replicate the masking distribution used in pre-training. We also replicate other pre-training practices, such as replacing 10% of masked tokens with random tokens and another 10% with the original tokens (devlin2018bert; lin2023evolutionary; su2023saprot) or cropping sequences to random 1024-token fragments (lin2023evolutionary; su2023saprot).
@@
Since customization with ProteinTTT does not assume more than a single protein, early stopping on validation data is not feasible. To address this, we first fine-tune the pre-trained parameters $\theta_{0}$ of a backbone $f$ for a fixed number of steps $T$, yielding parameters $\Theta=\{\theta_{0},\theta_{1},\dots,\theta_{T}\}$. The final customized parameters $\theta_{x}$ are selected as $\operatorname*{arg\,max}_{\theta\in\Theta}c(h(f(x;\theta)))$ where $c$ is a confidence function. If $c$ is not available, we set $\theta_{x}=\theta_{T}$. Section˜H.2 discusses how using pLDDT as the confidence function $c$ for structure prediction makes ProteinTTT robust to hyperparameter selection and how the number of steps $T$ can be fixed (e.g., $T=30$) while optimizing learning rate and batch size effectively. Before customizing for the next target protein, the parameters are reset to $\theta_{0}$.
To make ProteinTTT easily applicable to large-scale models (e.g., the 3B-parameter ESM2 backbone), we leverage low-rank adaptation (LoRA; hu2021lora) and gradient accumulation during customization. Additionally, to improve the stability and predictability of customization, we use stochastic gradient descent (SGD; ruder2016overview) instead of the commonly used Adam optimizer (kingma2015adam), following (gandelsman2022test). Further details are provided in Appendix˜F.
@@ 3.2 Inference on downstream tasks
Since Y-shaped architectures are prevalent in protein machine learning, ProteinTTT can be straightforwardly applied to numerous tasks. In this work, we consider three standard problems: protein structure, fitness, and function prediction, and apply our method to corresponding well-established models. For structure prediction, we apply ProteinTTT to ESMFold (Figure˜3, lin2023evolutionary, HelixFold-Single (fang2023method), DPLM2 Bit-based (hsieh2025elucidating), and ESM3 (hayes2024simulating); for fitness prediction, we use ESM2 (lin2023evolutionary), SaProt (su2023saprot), ProSST (li2024prosst), MSA Transformer (rao2021msa), and ProGen2 (nijkamp2023progen2); and for function prediction, we apply ProteinTTT to ESM-1v-based (meier2021language) EnzymeExplorer (samusevich2024discovery) and ESM-1b-based (rives2021biological) Light attention (stark2021light).
In all models we consider, $f$ is a Transformer encoder operating on protein tokens, and $g$ is a masked language modeling head mapping embeddings to amino acid types. The downstream head $h$, however, varies strongly by task. For structure prediction, $h$ is a structure predictor: AlphaFold2-inspired modules in ESMFold, HelixFold-Single and DPLM2 Bit-wise (jumper2021highly), and a VQ-VAE decoder in ESM3 (razavi2019generating). For fitness prediction, $h$ outputs a single score; all methods perform zero-shot inference using $h\circ f$ via log likelihoods from $g$, with $h$ acting as a simple, parameter-free adaptation of $g$. For function prediction, $h$ is a classifier: a random forest in EnzymeExplorer (samusevich2024discovery) and a light attention module in (stark2021light).
@@ 4.1 Protein structure prediction
Protein structure prediction is the task of predicting 3D atom coordinates from an amino acid sequence. It is arguably one of the best-established problems in computational biology (jumper2021highly).
@@
To evaluate the performance of ProteinTTT, we employ CAMEO, a standard benchmark for protein folding. We use the validation and test folds from lin2023evolutionary, focusing only on targets with low-confidence predictions from the base ESMFold, as determined by pLDDT and perplexity (Section˜F.1). We use the standard TM-score (zhang2004scoring) and LDDT (mariani2013lddt) metrics to evaluate global and local structure prediction quality, respectively.
As baseline methods, we use techniques alternative to ProteinTTT for improving the performance of the pre-trained base models. In particular, the ESMFold paper proposes randomly masking 15% of amino acids in a protein sequence before inference, allowing for sampling multiple protein structure predictions from the regression ESMFold model (lin2023evolutionary). For each sequence, we sample a number of predictions equal to the total number of ProteinTTT steps and refer to this baseline as ESMFold + MP (Masked Prediction). As a baseline for ESM3, we use chain-of-thought iterative decoding, referred to as ESM3 + CoT, proposed in the ESM3 paper (hayes2024simulating).
@@ Results.
Customization with ProteinTTT consistently improves the performance of all the tested methods, ESMFold, HelixFold-Single, and ESM3, outperforming the masked prediction (ESMFold + MP) and chain-of-thought (ESM3 + CoT) baselines, as shown in Table˜2. Among the 18 challenging CAMEO test proteins, ProteinTTT significantly improved the prediction of 7, 4, 5, and 6 structures from ESMFold, DPLM2 Bit-based, HelixFold-Single, and ESM3, respectively, while only moderately disrupting the prediction of 2, 1, 1, and 1 structures, respectively (Figure˜A7). Remarkably, ProteinTTT improves DPLM2 Bit-based despite the absence of a confidence function (no trained pLDDT head available) and despite the model being pretrained via discrete diffusion, while still using the same masked-modeling objective for customization as for the other methods.
Most notably, ProteinTTT enables accurate structure prediction for targets that are poorly predicted with the original models. For instance, Figure˜1 presents a strongly improved structure predicted using ESMFold + ProteinTTT for the target that was part of the CASP14 competition and shown as an unsuccessful case in the original ESMFold publication (lin2023evolutionary, Fig. 2E). Another example is shown in Figure˜3, where ProteinTTT refined the structure prediction from a low-quality prediction (TM-score =0.29=0.29) to a nearly perfectly folded protein (TM-score =0.92=0.92). Figure˜A4 shows that ESMFold + ProteinTTT maintains computational efficiency of ESMFold, being an order of magnitude faster than AlphaFold2. Figure˜A10 additionally demonstrates the robustness of ESM3 + ProteinTTT to the choice of hyperparameters.